OpenStudy (anonymous):

If h(x) = 2x + 1, find the value of x if h(x) = 11. 5 23 6 10

OpenStudy (anonymous):

@ivettef365 @Nurali @Loser66

OpenStudy (nurali):

h(x) = 2x + 1 11=2x + 1 2x=11-1 2x=10 x=10/2 x=____

OpenStudy (anonymous):

23?

OpenStudy (ivettef365):

sorry mistake h(x) = 11 so substitute the 11 as follows: 11 = 2x + 1 solve for x

OpenStudy (anonymous):

so 5?

OpenStudy (ivettef365):

that is 11 - 1 = 2x 10 = 2x x = 10/2 x = 5